# 4Université de Strasbourg, CNRS UMR 7550, Observatoire astronomique de Strasbourg, 67000 Strasbourg, France 5DIO, Observatoire de Paris, CNRS, PSL, 75014, Paris, France 6IDRIS, CNRS, F-91403 Orsay, France

### v1

@@ 4 Results
Figure 4: Evolution of the validation loss subparts (natural, Sect. A.7 in Paper I) during training for our three successive models. The first 20 iterations are omitted. Scores computed over the MAIN cube are indicated every 200 iterations after 400.
4.1 MAIN cube scoring
We present results from three successive bootstrap trainings, BASE, BT1, and BT2. We show the evolution of all natural loss subparts on our proxy validation dataset as a function of the iteration for these three trainings in Fig. 4. The score on the MAIN cube is computed periodically to confirm the absence of overtraining. We observe a loss plateau at the beginning of the training. This is well visible in the BASE model loss, but absent from BT1 and BT2 losses due to the use of pre-trained layers. This plateau corresponds to a regime in which the network is yet incapable of identifying the relevant signal for the task. Its length depends strongly on the method setup and backbone, with the worst-case scenario being an infinite plateau. The subspace of method parameters and architecture for which the network overcomes this state is small and difficult to find, highlighting the difficulty of the task. Selecting only the few hundred brightest sources as targets reduces the plateau’s length, but this comes at the cost of fainter sources never being detected, resulting in a lower score. This is one of the motivations for using a bootstrap training approach, which allows the selection function to be refined over successive trainings. For a given selection function, there is no clear correlation between the plateau’s length and the maximum achievable score when varying other parameters. Here, we optimized our method to reach the best score regardless of the plateau’s length. The short-term instability of both the loss and the score is caused by the low average source density, resulting in substantial variations in the typical source content between two iterations. For each training, we select the iteration that results in the highest score on the MAIN cube as our final model state. In principle, the best iteration should be determined based on an independent validation dataset distinct from both the training and final test datasets. However, as discussed in Sect. 3.3.1, splitting the small LDEV catalog would cause stronger issues, leading us to consider the MAIN catalog as our validation dataset to identify the best iteration.
Table 4: . SDC2 scores for source catalogs from different teams and methods.
@@ 4.1 MAIN cube scoring
Table 4 presents a detailed score result comparison of the catalogs produced by our three models and of teams that scored above 10 000 points in the original challenge. This table highlights the number of candidate detections, the matches based on the scorer criteria, the false positives, the purity, and the average source score. A small description of each team’s method and a detailed comparison of the produced catalogs are provided in the SDC2 summary paper (Hartley et al., 2023). Here, we summarize a few additional observations regarding the original challenge scores. ML methods were used for all or part of the pipeline of several teams, notably by the two top-scoring teams and three out of four teams with the lowest submitted score. This illustrates both the potential and the intrinsic difficulty of implementing such approaches for analyzing astronomical data. Also, most teams that applied denoising as a preprocessing failed to achieve a high score. This step likely removes signal from fainter sources, which are only detectable through their complex 3D signal distribution in a high noise environment. Although the first few layers of our backbone can act as a denoiser, the network learn to identify which signal should be compressed or preserved regarding the specific SDC2 task. Most teams using classical approaches relied on the same software, SOFIA (Serra et al., 2015; Westmeier et al., 2021). Team SOFIA notably achieved third place in the challenge. The second best-scoring team, FORSKA-Sweden, used a 3D U-net for source segmentation, but characterized them with SOFIA as well. Still, the remaining score difference between the classical SOFIA method and the two top-scoring approaches based on 3D CNNs highlights a specific strength of this type of model in extracting complex 3D patterns in high noise contexts (Appendix B). Interestingly, this contrasts with the results of Barkai et al. (2023), which performed a comparative study on a few HI detection methods and observed that SOFIA and the V-net 3D-CNN-based method performed similarly. This highlights again the importance of the architecture, method setup, and quality of the training data. Although some teams published a detailed version of their pipeline after the challenge, such as FORSKA-Sweden (Håkansson et al., 2023), there is currently no published post-challenge score comparable to ours. To facilitate future comparisons, all models and catalogs presented in the paper are made publicly available (Sect. 6).
@@
Our new BASE model improves our top challenge score by $6.1\%$, and the BT1 model pushes this improvement to $9.5\%$, placing it $13.2\%$ higher than the second top score. The number of matched sources also increases by $10.6\%$ between our BT1 model and our top challenge score. We note that the typical variation in best score over multiple training of the same setup is typically around $\pm 0.5\%$. Additional bootstrap trainings after the first one produce no significant improvement, except for faster convergence, as illustrated by our BT2 model. This indicates that our base selection function is already a good starting point. In Appendix D, we show that the bootstrap approach can help reach similar scores when starting from a degraded selection function. We highlight that our MINERVA v0.1 score was achieved through a combination of two methods, a prototype version of YOLO-CIANNA, and a hybrid classical and ML method called CHADHOC (Sect. 4.1.7 from Hartley et al., 2023). At the time, the YOLO-CIANNA prototype was limited to $\sim$21 000 points, meaning that we improved our method-specific score by around $20\%$. The distribution of predicted sources is mostly homogeneous over all three dimensions of the MAIN cube. We represent cutouts centered on a few matched sources in Appendix B.
@@
Figure 5: (a) Two-dimensional histogram of the MAIN cube detection matches target volume as a function of their target integrated line flux. The red dashed line represents the selection function as in Fig. 2. (b) Histogram of the target line flux for the MAIN truth catalog and the matched sources. False positives are added based on their predicted flux. Completeness is overplotted for relevant bins. (c) Histogram of the predicted line flux for the matched predicted sources and false positives. The MAIN truth catalog is plotted in the background based on the target line flux. Purity is overplotted for relevant bins. (d) Histogram of the S/N for the MAIN truth catalog and the matched sources. Completeness is overplotted for relevant bins.
@@
We represent histograms of the integrated line flux for both the matched and false detections in comparison to the complete MAIN truth catalog for BT1 in Fig. 5. The first one is based on the target values and used to compute the completeness as a function of the real line flux, whereas the second one is based on the predictions and used to calculate the purity as a function of the predicted line flux. We also represent a histogram of the matched prediction smoothed S/N, as defined in Sect. 6.2 of Hartley et al. (2023). This S/N estimate was added to the MAIN truth catalog after the challenge and is therefore only used for analyzing the results. Finally, this figure shows a 2D histogram of the matched sources volume against the target line flux. These plots show that our method can generalize outside our selection function, which is reinforced by the bootstrap steps, but is also the case for our BASE model. False detections arise mainly in the low flux regimes where completeness and purity drop smoothly.
@@
Our method produces a list of detections ordered by objectness confidence. Instead of selecting the threshold that maximises the SDC2 score, we can search for a threshold that enforces a purity of at least $99\%$. We report the scores and properties of these purity-optimized catalogs for our three models in Table 4. Interestingly, the resulting catalogs still score 10 to $17\%$ higher than team SOFIA, the best non-ML approach. On the contrary, lowering the threshold results in more detection candidates. When applied to real observational data, these low-confidence detections could serve as follow-up targets in an observing program that aims to confirm some of them. The real strength of predicting a self-assessed confidence score is that a trained model samples the full detectability range. Therefore, as we did in Sect. 5.1.4 of Paper I, we compute an alternative metric similar to an average precision (AP). For this, we lower our initial objectness prediction clip to 0.05, order the prediction by objectness, and pass it to the SDC2 scorer to obtain the list of matches. We then compute the running precision (purity) and recall (completeness) over the catalog and convert them into a smoothed precision-recall curve, whose integral defines the SDC2-based AP. We report AP values of 18.35, 18.32, and 18.34 for our BASE, BT1, and BT2 models, respectively. We also find the number of matches found regardless of the objectness value to be around 50 000 for all three models. To estimate the chances of random matches with a background source, we compute the AP of the BASE catalog with the 3D positions randomized, which produces only 361 false matches. This indicates that detections in the low objectness regime contain a fair amount of real detections, but at a purity lower than the SDC2 scorer requirement. We also compute a purely geometrical DIoU-based AP score for all models, which is presented in Appendix C.
@@ 4.2 Source characterization quality
Figure 6: Subpart error distributions, presented as 2D histograms against relevant target quantities or scales, along with the corresponding 1D error histogram and the associated score response functions. The red dashed lines indicate the error region for which the score is saturated to one. Orange lines and arrows show the range of values from the training sample when relevant. Dotted black lines represent the error obtained when predicting the minimum value from the training sample across the full comparison scale.
@@
Our BT1 model achieves an average characterization score of $\bar{s}=0.8298$, improving our original challenge submission by 0.02, which was already the highest characterization score among all teams. Multiplied by the approximate 34 000 matches, this results in a direct increase of about 680 points. However, an average characterization score is not suitable for comparing methods with varying detection performances. This penalizes strong detectors that successfully detect many faint sources, which are likely also difficult to characterize. This is illustrated by our alternative purity-optimized catalogs that all achieve $\bar{s}>0.85$ by selecting only the most obvious detections. A better approach would be to evaluate the average score on the catalog intersections for each pair of methods to be compared, so it is computed on the same list of sources. However, to date, the catalogs submitted by teams that participated in the original challenge have not been publicly distributed. This is one of the motivations for distributing all our SDC2 catalogs along with the paper, which should enable future comparison (Sect. 5).
@@
We also highlight that the scorer uses relative errors for which the positive and negative sides are asymmetric, and for which the positive side is unbound and the negative side is limited to minus one. Combined with symmetric score response functions, it implies that catalogs that systematically underestimate the parameters achieve higher scores than those with a zero-centered error distribution. This limit is inherited from the SDC1 scorer as discussed in Sect. 5.1.3 of Paper I.
@@
To thoroughly evaluate the characterization accuracy of our method, we rely on the detailed scoring information exposed by the scorer code. For each match, it provides the predicted and target source properties, the matching distance, and all error subpart values $E$ as presented in Table 1, as well as the final source score recombined by Eq. 7. From this, we can draw individual error subparts histograms and compare the resulting error distribution to the associated score response function, which are represented in grey in Fig. 6. For each error subpart, we indicate the region of the distribution where the score is saturated using red dashed lines. The more the error distribution is contained in this area, the higher the average subpart score. To push the analysis further, for each error subpart, we identify an appropriate source target property against which it would be interesting to represent the error distribution. For example, we represent the flux error as a function of the target’s integrated line flux, which is natural considering that $E^{\rm flux}$ is simply a relative flux error. The sky position, the HI size, and PA are represented against the target beam-convolved HI size. The frequency position, the line width, and the inclination are represented against the target line width. All these distributions are represented in the form of 2D histograms, presented alongside their 1D counterparts in Fig. 6.
@@
From this figure, we observe that faint sources exhibit a strong positive error. This results from multiple effects. Firstly, the cube noise imposes a hard limit on the minimum predictable flux value. Secondly, the random fluctuation of the noise can either boost or reduce the apparent flux. Therefore, sources with a per-voxel flux close to the noise level will only be detected when the noise contribution is positive, acting as a selection effect of sources with overestimated flux. This effect is reinforced by the detector being unlikely to predict a flux lower than the minimum value it has seen during training. This is illustrated by the dashed black line over the 2D $E^{\rm flux}$ error distribution representing the error that would arise by systematically predicting the minimum training flux. Considering that the scorer includes the flux error in its matching criteria (Eq. 1), properly detected faint sources might be rejected based on high flux error, even though we reached the best achievable flux prediction for the noise level. Although the scorer does not expose the $D_{\rm tot}$ value for rejected sources, we observe that the faintest matches already have values close to the $D_{\rm tot}\geq 5$ rejection limit, suggesting that some of our false positives might fall in this category. On the contrary, the flux of the brightest sources is underestimated, which is mainly due to the scarcity of bright examples in the training sample.
@@
We observe similar behavior for both the source size and line width errors, also resulting from observational limits, specifically the beam size and channel resolution. These subparts have high average scores due to less strict error thresholds. Both the sky position and the central frequency are well characterized, which arises from the fact that poorly positioned detections have already been rejected by the matching criteria that are more strict than the associated subscores. We note that despite their resemblance, $E^{\rm pos}$ and $D_{\rm pos}$ are two different quantities, as they rely on different reference sizes ($S^{\prime}\neq S^{\prime\prime}$), which explains why the position error distribution does not expand up to the matching rejection limit of one. Finally, the position angle is well characterized for sources with a sufficient HI size but predicted at random for small sources, which is also the result of the observational limit imposed by the beam size. The inclination angle achieves a better score. However, the substructures visible in the distribution as a function of the line width suggest that this angle might be inferred from correlated parameters rather than directly measured from the cube. We confirm that the network learns parameter correlations to some degree in Appendix A by representing the source distribution for a selected set of parameter couples.

### v2

@@ 4 Results
Table 5: SDC2 scores for source catalogs from different teams and methods.
@@
4.1 MAIN cube scoring
We present results from four successive bootstrap trainings: BASE, BT1, BT2, and BT3. We show the evolution of all natural loss subparts on our proxy validation dataset as a function of the iteration for these four trainings in Fig. 4. We also represent the evolution of the score evaluated on the MAIN cube as a function of the iteration to confirm the absence of overtraining. We observe a plateau in the loss at the beginning of training for the BASE model. This plateau corresponds to a regime in which the network is still unable to identify the relevant signal for the task. Its length depends strongly on the method setup and backbone, with the worst-case scenario being an infinite plateau. The subspace of method parameters and architecture that enables the network to overcome this state is narrow and hard to find, highlighting the difficulty of the task. We observed that selecting only the few hundred brightest sources as targets reduces the plateau’s length, but it also means fainter sources are never detected, resulting in a lower score. This is one of the motivations for using a bootstrap training approach, which allows the selection function to be refined over successive trainings. For a given selection function, there is no clear correlation between the plateau’s length and the maximum achievable score when varying other parameters. Here, we optimized our method to reach the best score regardless of the plateau’s length. BT1 to BT3 training losses do not exhibit the same initial loss plateau because these models rely on pre-trained layers from their respective previous bootstrap models, which had already learned the task’s principles. We discuss how bootstrapping using from-scratch retraining of each model affects this initial plateau in Sect. D. The observed short-term instability in both loss and score is caused by the low average source density, which leads to strong variation in typical source content between iterations. For each training, we selected the iteration with the highest score on the MAIN cube as the final model state. In principle, the best iteration should be determined based on an independent validation dataset distinct from both the training and final test datasets. However, as discussed in Sect. 3.4, splitting the small LDEV catalog would further reduce an already small training source catalog.
@@ 4.1 MAIN cube scoring
Table 5 presents a detailed score result comparison of the catalogs produced by our four models and of teams that scored above 10 000 points in the original challenge. This table highlights the number of candidate detections, matches based on the scorer criteria, false positives, purity, and the average source score. A small description of each team’s method and a detailed comparison of the produced catalogs are provided in the SDC2 summary paper (Hartley et al., 2023). Here, we summarize a few additional observations regarding the original challenge scores. ML methods were used for all or part of the pipeline of several teams, notably by the two top-scoring teams and three of the four teams with the lowest submitted scores. This illustrates both the potential and the intrinsic difficulty of implementing such approaches for analyzing astronomical data. Also, most teams that applied denoising as a preprocessing failed to achieve a high score. This step likely removes signal from fainter sources, which are detectable only through their complex 3D signal distribution in a high noise environment. Although the first few layers of our backbone can act as a denoiser, the network learns which signal to compress or preserve regarding the specific SDC2 task. Most teams using classical approaches relied on the same software, SoFiA (Serra et al., 2015; Westmeier et al., 2021). Team SOFIA notably achieved third place in the challenge. The second best-scoring team, FORSKA-Sweden, used a 3D U-Net for source segmentation but also relied on SoFiA for a part of its source characterization. Still, the remaining score difference between the classical SoFiA method and the two top-scoring 3D-CNN-based approaches highlights a specific strength of this type of model in extracting complex 3D patterns in high noise contexts (Appendix B). Interestingly, this contrasts with the results of Barkai et al. (2023), which performed a comparative study of a few HI detection methods and observed that SoFiA and a V-net 3D-CNN-based method performed similarly. This also highlights the importance of the architecture, method setup, and quality of the training data. Although some teams published a detailed version of their pipeline after the challenge, such as FORSKA-Sweden (Håkansson et al., 2023), there is currently no published post-challenge score comparable to ours. To facilitate future comparisons, all models and catalogs presented in this paper are made publicly available (Sect. 6).
@@
Our new BASE model improves our top challenge score by $6.1\%$, and the BT1 model pushes this improvement to $9.5\%$, placing it $13.2\%$ higher than the second top score. The number of matched sources also increases by $10.6\%$ between our BT1 model and our top challenge score. We note that the typical variation in best achievable score over multiple training using the same setup is typically around $\pm 0.6\%$. Additional bootstrap training steps after the first one produce no significant improvement, except for faster convergence, as illustrated by our BT2 and BT3 models. The BT1 model scoring higher is mostly due to chance, as the training sample for all three BT models does not change much (see Sect. 3.5). BT2 and BT3 scores are within typical retraining variability using the same setup. This indicates that our base selection function is already a good starting point. In Appendix E, we show that the bootstrap approach can help reach similar scores when starting from a degraded selection function. We highlight that our MINERVA v0.1 score was achieved through a combination of two methods, a prototype version of YOLO-CIANNA, and a hybrid classical and ML method called CHADHOC (Sect. 4.1.7 from Hartley et al., 2023). At the time, the YOLO-CIANNA prototype was limited to $\sim$21 000 points, meaning that we improved our method-specific score by around $20\%$. The distribution of predicted sources is mostly homogeneous over all three dimensions of the MAIN cube. We represent cutouts centered on a few matched sources in Appendix B.
@@
Figure 5: (a) Two-dimensional histogram of the MAIN cube detection matches target volume as a function of their target integrated line flux. The dashed red line represents the selection function as in Fig. 2. (b) Histogram of the target line flux for the MAIN truth catalog and the matched sources. False positives are added based on their predicted flux. Completeness is overplotted for relevant bins. (c) Histogram of the predicted line flux for the matched predicted sources and false positives. The MAIN truth catalog is plotted in the background based on the target line flux. Purity is overplotted for relevant bins. (d) Histogram of the smoothed S/N for the MAIN truth catalog and the matched sources. Completeness is overplotted for relevant bins. (e) Identical to d but using the raw non-smoothed S/N.
@@
We represent histograms of the integrated line flux for both the matched and false detections in comparison to the complete MAIN truth catalog for BT1 in Fig. 5. The first histogram is based on the target values and used to compute completeness as a function of the real line flux, whereas the second is based on the model predictions and used to calculate purity as a function of the predicted line flux. We also represent histograms of the target smoothed and non-smoothed S/N for the matched detections, as defined in Sect. 6.2 of Hartley et al. (2023). This S/N estimate was added to the MAIN truth catalog after the challenge and is therefore only used for analyzing the results. Finally, this figure shows a 2D histogram of the matched source volume against the target line flux. These plots show that false detections arise mainly in the low-flux regimes where completeness and purity drop smoothly. We also observe that the model correctly detects sources fainter than the limit imposed by our selection function, which is primarily due to our bootstrapping approach. Still, this ability to detect sources outside the training distribution is also observed in our BASE model’s predictions, though to a lesser extent. These faint sources are likely rendered detectable by flux boosting from the local noise, as discussed in the next section.
@@
Our method produces a list of detections ordered by objectness confidence. Instead of selecting the threshold that maximizes the SDC2 score, we can search for a threshold that enforces a purity of at least $99\%$. We report the scores and properties of these purity-optimized catalogs for our four models in Table 5. The resulting catalogs still score 10 to $17\%$ higher than team SOFIA, the best non-ML approach. On the contrary, lowering the threshold results in more detection candidates. When applied to real observational data, these low-confidence detections could serve as follow-up targets for an observing program aimed at confirming some of them. The advantage of predicting a self-assessed confidence score is that a trained model samples the full detectability range. Therefore, as we did in Sect. 5.1.4 of Paper I, we computed an alternative metric similar to an average precision (AP). For this, we lowered our initial objectness prediction threshold to 0.05, sorted the predictions by objectness, and passed them to the SDC2 scorer to obtain the list of matches. We then computed the running precision (purity) and recall (completeness) over the catalog and converted them into a smoothed precision-recall curve, whose integral defines the SDC2-based AP. We report AP values of 18.35, 18.32, 18.34, and 18.25 for our BASE, BT1, BT2, and BT3 models, respectively. The total number of matches is close to 50 000 for all models. To estimate the chances of random matches with a background source, we computed the AP of the BASE catalog after randomizing the 3D positions of all detections, yielding only 361 fortuitous matches. This indicates that detections in the low objectness regime contain a fair number of real detections, but the number of false detections is likely too high relative to the SDC2 scorer requirement. We also computed a purely geometric DIoU-based AP score for all models, which is presented in Appendix C.
@@ 4.2 Source characterization quality
Figure 6: Subpart error distributions as 2D histograms against relevant target quantities or scales. For each subpart, the corresponding 1D error histogram and the associated score response functions are provided. The dashed red lines indicate the error region where the score saturates at 1. Orange lines and arrows indicate the range of values from the training sample when relevant. Dotted black lines represent the error obtained when predicting the minimum value from the training sample across the full comparison scale.
@@
Our BT1 model achieves an average characterization score of $\bar{s}=0.8298$, improving our original challenge submission by 0.02, which was already the highest characterization score among all teams. When multiplied by the approximate 34 000 matches, this results in a direct increase of about 680 points. However, an average characterization score is not suitable for comparing methods with varying detection performances. This penalizes strong detectors that successfully detect many faint sources, which are likely also difficult to characterize. This is illustrated by our alternative purity-optimized catalogs that all achieve $\bar{s}>0.85$ by selecting only the most obvious detections. A better approach would be to evaluate the average characterization score for catalog intersections for each pair of methods to compare, so it is computed on the same list of sources. However, to date, the catalogs submitted by teams that participated in the original challenge have not been publicly distributed. This is one of the motivations for distributing all our SDC2 catalogs, which should enable future comparisons (Sect. 5).
@@
We also note that the scorer uses relative errors, which lead to an asymmetry between positive and negative errors, with the positive side unbounded and the negative side limited to minus one. When combined with symmetric score response functions, catalogs that systematically underestimate parameters achieve higher scores than those with a zero-centered error distribution. This bias is inherited from the SDC1 scorer design as discussed in Sect. 5.1.3 of Paper I.
@@
To thoroughly evaluate the characterization accuracy of our method, we rely on the detailed information exposed by the scorer code. For each match, it provides the predicted and target source properties, the matching distance, all error subpart values $E$ from Table 1, and the final source score recombined by Eq. 7. From this, we draw individual error subpart histograms and compare the resulting error distribution with the associated score response function, both shown in gray in Fig. 6. For each error subpart, we use dashed red lines to indicate the region of the distribution where the score is saturated. The more the error distribution is contained in the corresponding area, the higher the average subpart score. For each error subpart, we identified an appropriate source target property against which it is relevant to represent the error distribution. For example, we represent the flux error as a function of the target’s integrated line flux, which is natural considering that $E^{\rm flux}$ is simply a relative flux error. The sky position, HI size, and PA are represented against the target beam-convolved HI size. The frequency position, line width, and inclination are represented against the target line width. All these distributions are represented in the form of 2D histograms, presented alongside their 1D counterparts in Fig. 6.
@@
From this figure, we observe that faint sources exhibit a strong positive error. This results from multiple effects. Firstly, the cube noise imposes a hard limit on the minimum predictable flux. Secondly, the random fluctuation of the noise can either boost or reduce the apparent flux. Sources with per-voxel flux near the noise level can only be detected when the noise contribution is positive, thereby acting as a selection effect of sources with overestimated flux. This effect is reinforced by the detector’s inability to predict a flux below the minimum target value seen during training. This is illustrated by the dashed black line over the 2D $E^{\rm flux}$ error distribution representing the error that would arise by systematically predicting the minimum training flux. Considering that the scorer includes the flux error in its matching criteria (Eq. 1), properly detected faint sources might be rejected based on high flux error, even though we reached the best achievable flux prediction for the noise level. Although the scorer does not expose the $D_{\rm tot}$ value for rejected sources, we observe that the faintest matches already have values close to the $D_{\rm tot}\geq 5$ rejection limit, suggesting that some of our false positives might fall in this category. On the contrary, the flux of the brightest sources is underestimated, which is mainly caused by the scarcity of bright examples in the training sample.
@@
We observe similar behavior in both source size and line width error distributions, which also result from observational limits, specifically the beam size and channel resolution. These subparts have higher average scores due to less strict error thresholds. Both the sky position and the central frequency appear well characterized, as would likely be the case for any catalog of matches, since poorly positioned detections have already been rejected by matching criteria stricter than the associated subscores. We note that despite their resemblance, $E^{\rm pos}$ and $D_{\rm pos}$ are two different quantities, as they rely on different reference sizes ($S^{\prime}\neq S^{\prime\prime}$), which explains why the position error distribution does not expand up to the matching rejection limit of one. Finally, the position angle is well characterized for sources with sufficient HI size but is predicted at random for small sources, which is also a consequence of the observational limit imposed by the beam size. The inclination angle achieves a better score. However, the substructures visible in the distribution as a function of the line width suggest that this angle might be inferred from correlated parameters rather than directly measured from the cube. We confirm that the network learns parameter correlations to some degree in Appendix A by representing the source distribution for a selected set of parameter couples.